OpenStudy (anonymous):

Solve the following equation to get y by itself: 4y + 2z = 20

jimthompson5910 (jim_thompson5910):

To isolate y, we first subtract 2z from both sides \[\Large 4y + 2z = 20\] \[\Large 4y + 2z{\color{red}{-2z}} = 20{\color{red}{-2z}}\] \[\Large 4y + 0z = 20-2z\] \[\Large 4y + 0 = 20-2z\] \[\Large 4y = 20-2z\] ------------------------------------------ Then we divide both sides by 4 to fully isolate y \[\Large 4y = 20-2z\] \[\Large \frac{4y}{{\color{blue}{4}}} = \frac{20-2z}{{\color{blue}{4}}}\] \[\Large y = \frac{20-2z}{4}\] \[\Large y = \frac{-2z+20}{4}\] \[\Large y = \frac{-2(z-10)}{2*2}\] \[\Large y = \frac{-\cancel{2}(z-10)}{\cancel{2}*2}\] \[\Large y = -\frac{z-10}{2}\]